an auditorium has 50 rows of seats. THe first row has 20 seats, the row has 21 seats, the third row has 22 seats, and so on. each row having one more seat than the previous row. How many seats are there all together?
Answer #1
First row = 20 seats,
second row = 21 seats,
...
nth row = 19+n seats,
...
50th row = 19+50 = 69 seats

Write the number of seats in forward and reverse order:
19 + 69 = 88
20 + 68 = 88
...
44 + 44 = 88
...
69 + 19 = 88
---------
Sum = 50*(88/2) = 2200 seats
answered by: score
Answer #2
19+n
u#1=19+1
u#1=20
u#50=19+50
u#1=69
50(20+69)/2
answered by: Laury